Mindy wrote down a decimal number. The digit in the tenth's place is half the digit in the hundredth's place. If the product of the two digits is 18, what is the number?

Answer by checkley77(12844)

LET X BE THE TENS POSITION DIGIT & Y THE HUNDRETHS DIGIT.
XY=18 OR X=18/Y
Y/2=X
Y/2=18/Y CROSS MULTIPLY.
Y^2=36
Y=SQRT36
Y=6 ANS.
6/2=X
X=3 ANS.
.36 IS THE ANSWER.
PROOF
6*3=18